OpenStudy (anonymous):

what is cos(theta/2) if tan(theta)=8/15

OpenStudy (richyw):

\[\tan\theta=\frac{8}{15}\]\[\theta=\tan^{-1}\left(\frac{8}{15}\right)\]\[\cos\left(\frac{\theta}{2}\right)=\cos\left(\frac{\tan^{-1}\left(\frac{8}{15}\right)}{2}\right)\]

OpenStudy (richyw):

\[=\frac{4}{\sqrt{17}}\]